a) The force of gravity does 405.7 J of work on an 18 kg object as it is lifted vertically with a   constant velocity. How high was the object lifted? 

b) What type of work is done by the force of gravity as the object was lifted? Explain.  

 

 

 

 

 

 



1
Expert's answer
2021-06-02T10:04:29-0400

(a)

"W_g=mgh,""h=\dfrac{W_g}{mg}=\dfrac{405.7\ J}{18\ kg\cdot9.8\ \dfrac{m}{s^2}}=2.3\ m."

(b) The force of gravity does negative work as the object was lifted. The work done is negative because displacement directed in the opposite direction to the force of gravity.